# C-MORL: Multi-Objective Reinforcement Learning through Efficient Discovery of Pareto Front

### v1

@@ Abstract
Multi-objective reinforcement learning (MORL) excels at handling rapidly changing preferences in tasks that involve multiple criteria, even for unseen preferences. However, previous dominating MORL methods typically generate a fixed policy set or preference-conditioned policy through multiple training iterations exclusively for sampled preference vectors, and cannot ensure the efficient discovery of the Pareto front. Furthermore, integrating preferences into the input of policy or value functions presents scalability challenges, in particular as the dimension of the state and preference space grow, which can complicate the learning process and hinder the algorithm’s performance on more complex tasks. To address these issues, we propose a two-stage Pareto front discovery algorithm called Constrained MORL (C-MORL), which serves as a seamless bridge between constrained policy optimization and MORL. Concretely, a set of policies is trained in parallel in the initialization stage, with each optimized towards its individual preference over the multiple objectives. Then, to fill the remaining vacancies in the Pareto front, the constrained optimization steps are employed to maximize one objective while constraining the other objectives to exceed a predefined threshold. Empirically, compared to recent advancements in MORL methods, our algorithm achieves more consistent and superior performances in terms of hypervolume, expected utility, and sparsity on both discrete and continuous control tasks, especially with numerous objectives (up to nine objectives in our experiments).
@@ 1 Introduction
In many real-world control and planning problems, multiple and sometimes even conflicting objectives are getting involved. Such situations necessitate striking a better trade-off among these decision-making goals (Roijers et al., 2013; Hayes et al., 2022). For instance, in industrial control scenarios (Salvendy, 2001; Wang et al., 2023), maximizing utility and minimizing energy consumption are of particular interest as objectives to be optimized. Since different decision makers have heterogeneous preferences over these objectives, there may exist multiple Pareto-optimal policies (Roijers et al., 2014). Classical reinforcement learning (RL) methods typically involve training individual policies exclusively to align with each preference weight vector over multiple rewards (Nagabandi et al., 2018; Gupta et al., 2018). Yet it may lead to an enormous computational burden due to the overly dependence on the model retraining and fine-tuning stages. Moreover, such policies are hard to directly generalize or transfer to newer tasks (Cobbe et al., 2019; Taiga et al., 2022). Therefore, the multi-objective reinforcement learning (MORL) paradigm has drawn significant attention by reformulating these tasks for optimizing towards multiple criterion (Xu et al., 2020; Basaklar et al., 2022; Zhu et al., 2023). MORL aims to obtain either a single policy readily adapted to different preferences (Felten et al., 2023b; 2022; Teh et al., 2017) or a set of policies (Zhao and Grover, 2024; Felten et al., 2024; Röpke et al., 2024; Kim et al., 2024) aligned with their respective preferences.
@@
Our algorithm achieves favorable time complexity characteristics and derives a high-quality Pareto front, as indicated by the following observations. Firstly, training several initial policies based on fixed preferences is efficient. In addition, the adoption of constrained update steps allows for rapid adjustment of the initial policy, leading to the derivation of new solutions on the Pareto front. Regarding the second research gap concerning the complete Pareto front, different from the meta-policy approach that relies on a single initial policy (Chen et al., 2019), our method can extend multiple policies selected from initial policies based on their crowd distance to enhance diversity while promoting better performances. Intuitively, a larger crowd distance indicates that the corresponding policy appears on a sparser area on the Pareto front, therefore, extending such a policy is more likely to fill the Pareto front. While sharing similarities with our formulation, as one category of multi-objective optimization (MOO) approach, epsilon-constraint methods solves a MOO problem by converting it to several single-objective constrained optimization problems (Laumanns et al., 2006; Van Moffaert and Nowé, 2014). However, the running time of such a method is exponential in problem size, rendering it impractical for MOO problems with numerous objectives. In contrast, the crowd-distance-based policy selection eases such burden, exhibiting linear complexity, and therefore can efficiently solve MORL tasks with numerous objectives.
In summary, our main contributions are as follows:
@@ 2 Related Work
Prior trials on tackling multi-objective RL fall into training single preference-conditioned policy or multi-policy. Typical single-preference-conditioned policy approaches adopt a policy that takes preferences as part of network inputs and utilizes the weighted-sum scalarization of the value functions (or advantage functions) to optimize the policy  (Van Moffaert et al., 2013; Parisi et al., 2016; Yang et al., 2019; Zhang and Golovin, 2020; Basaklar et al., 2022; Lu et al., 2022; Hung et al., 2022; Zhu et al., 2023; Lin et al., 2024). In the evaluation stage, agents can execute corresponding solutions based on users’ desired preferences (Yang et al., 2019; Basaklar et al., 2022). With the shared neural networks, gradients from different tasks can interfere negatively, making learning unstable, leading to imbalanced performance across the entire preference space, and sometimes even less data efficient.
@@
Instead of training a single policy agent, multi-policy approaches train a finite set of policies to approximate the Pareto front (Abels et al., 2019; Xu et al., 2020; Alegre et al., 2022). In the evaluation stage, the policy in policy set that maximizes the utility of the user (i.e., the weighted sum of objectives) is chosen (Friedman and Fontaine, 2018). Rather than evenly sampling from preference space, (Xu et al., 2020) proposes an efficient evolutionary learning algorithm to find the Pareto set approximation along with a prediction model for forecasting the expected improvement along vector-valued objectives. Yet the performance highly depends on prediction model’s accuracy, and it is challenging to recover the Pareto front due to the long-term local minima issue. This method also suffers from low training efficiency due to the exponential complexity involved in repeatedly calculating the quality of the virtual Pareto front during the training process. (Kyriakis and Deshmukh, 2022) designs a policy gradient solver to search for a direction that is simultaneously an ascent direction for all objectives. (Alegre et al., 2022) firstly trains a set of policies whose successor features form a $\epsilon$-$CCS$ (convex coverage set), then utilizes the generalized policy improvement (GPI) algorithm to derive a solution for a new preference. However, the training of the CCS policy cannot be parallelized. (Alegre et al., 2023) further introduces a novel Dyna-style MORL method that significantly improves sample efficiency while relying on accurate learning of dynamics.
@@
Our approach has connections with (Chen et al., 2019) and (Huang et al., 2022), while distinguishing itself in terms of the training process and objectives. Compared to the meta-policy approach, adaptation process is not necessary in the evaluation phase of our method. When presented with an unseen preference, the policy in the Pareto front with the highest utility (i.e., with the largest projection length of its return along the preference vector) is chosen as the surrogate execution policy. In contrast to (Huang et al., 2022; Kim et al., 2024), we do not aim at solving a constrained RL problem for training the working policy under specific preference. Rather, we leverage the constrained optimization steps to fill the complete Pareto front with enhanced flexibility (Liu et al., 2020; Xu et al., 2021), and design extension and selection algorithms to explicitly promote diverse policies. Constrained optimization techniques are widely adopted to solve multi-objective optimization (MOO) problems. Epsilon-constraint methods are a category of MOO techniques that involve pre-defining a virtual grid in the objective space and solving single-objective problems for each grid cell, where the optimum of each problem corresponds to a Pareto-optimal solution (Laumanns et al., 2006). Instead of pre-defining virtual grid, this work proposes crowd distance based policy selection to address the exponential complexity in the epsilon-constraint method.
@@ 3.1 MORL Setup
The goal of MORL is to find a policy so that each objective’s expected return in $\bm{G}^{\pi}$ can be optimized. In practice, since training RL typically requires a scalar reward to interact with the training agent, and co-optimizing multiple objectives is hard to achieve an ideal tradeoff, especially in a situation where objectives are competing against each other. To that end, denote $\Omega=\{\bm{\omega}\in\Omega|\sum_{i=1}^{n}\omega_{i}=1\}$ as the preference vector. We use the preference function $f_{\bm{\omega}}(\mathbf{r})$ to map a reward vector $\mathbf{r}(\mathbf{s},\mathbf{a})$ to a scalar utility given $\bm{\omega}:f_{\bm{\omega}}(\mathbf{r}(\mathbf{s},\mathbf{a}))=\bm{\omega}^{\top}\mathbf{r}(\mathbf{s},\mathbf{a})$. Our goal is then to find a multi-objective policy $\pi(\mathbf{a}|\mathbf{s},\bm{\omega})$ such that the expected scalarized return $\bm{\omega}^{\top}\bm{G}^{\pi}$ is maximized.
@@ 3.2 Pareto Optimality
Figure 1: Visualization of (a) hypervolume, reference point, and example of crowd distance calculation; and (b) expected utility calculated by assigning the set max policy from Pareto front solutions with the highest utility.
@@ 4 Optimizing Constrained MORL
Assumption 4.2 justifies the use of gradient-based algorithm for solving CMDP can converge with general constraints. For the MORL problem, as long as the threshold $\mathbf{d}$ are chosen beforehand such that for the initial points, there are $G_{i}^{\pi}\geq d_{i},i\neq l$, the initialization point of problem Eq. 1 is always feasible.
@@
$D^{*}\triangleq\min_{\bm{\lambda}}\max_{\pi}\;L(\bm{\lambda},\pi)$ (2)
@@ 5.3 Pareto Extension
where $G_{i}^{\pi_{r}}$ is the expected return of the $i^{th}$ objective in last constrained optimization step, which is indexed by $r$ for the current iteration.
C-MORL-CPO Rather than using sampling-based approaches (Duan et al., 2016) for finding policy updates in the relaxed formulation, empirically we follow the adapted trust region method (Schulman et al., 2015) for CPO updates (Achiam et al., 2017):
C-MORL-IPO Although CPO algorithm monotonically improves the policy and guarantees constraint satisfaction during the training process, an inner-loop optimization is needed when there are multiple constraints. Therefore, CPO is hard to handle more than two MORL objectives. To overcome this issue, inspired by (Xu et al., 2021; Liu et al., 2020), we also propose to solve C-MORL via interior point method (IPO), which is a primal type approach and holds the promise of finding solution closer to original C-MORL as shown in Theorem 4.5. Given $G_{i}^{\pi_{r}}$ and the the defined log barrier in 4.5, we then convert problem Eq. 1 into an unconstrained optimization problem:
@@
The detailed procedures of solving Eq. equation 4 and Eq.equation 5 are provided in Appendix F. In practice, IPO is more robust in stochastic environments, and larger $t$ would guide to a solution with higher rewards yet with more computation costs. Without the requirement of calculating the second-order derivative, C-MORL-IPO is more computationally efficient than C-MORL-CPO.
@@ 6 Experiments
In this Section, we validate the design of our proposed algorithm using both popular discrete and continuous MORL benchmarks from MO-Gymnasium (Felten et al., 2023a) and SustainGym (Yeh et al., 2024). These benchmarks include five comprehensive domains: (i) Grid World includes Fruit-Tree, a discrete benchmark with six objectives. (ii) Classic Control includes MO-Lunar-Lander, a discrete benchmark with four objectives. (iii) Miscellaneous includes Minecart, a discrete benchmark with four objectives. (iv) Robotics Control includes five MuJoCo tasks with continuous action space based on MuJoCo simulator (Todorov et al., 2012; Xu et al., 2020). (v) Sustainable Energy Systems includes two building heating supply tasks. These benchmarks pose significant challenges for MORL due to complex state (up to $376$)/action (up to $23$) spaces and large objectives spaces (up to $9$ objectives). More details of experiment settings are listed in Appendix E.
We benchmark our algorithm against five competitive baselines under the metrics of hypervolume (HV), expected utility (EU), and sparsity metrics (SP). Higher HV and EU, lower SP indicate better performance. Each of the baselines are trained for $5\times 10^{5}$ time steps for discrete benchmarks. Continuous benchmarks with two, three, and nine objectives are trained for $1.5\times 10^{6}$, $2\times 10^{6}$, and $2.5\times 10^{6}$ steps, respectively. The baselines involve: (i)Envelope (Yang et al., 2019).(ii) CAPQL (Lu et al., 2022). (iii)Q-Pensieve (Hung et al., 2022). (iv) PG-MORL (Xu et al., 2020). (v) GPI-LS (Alegre et al., 2023). Among these baselines, Envelope is developed for discrete control tasks, GPI-LS is suitable for both discrete and continuous control. The other baselines are specifically tailored for continuous control.
@@
While C-MORL leads the highest hypervolume by a large margin in almost all cases, it does not always attain the best sparsity in some benchmarks. This occurs because, in certain cases, an algorithm may identify only a few similar Pareto-optimal solutions, leading to low sparsity. As illustrated in Table 2, along with Figures 4 (c) and 5, both CAPQL and GPI-LS cover only specific portion of the Pareto front, resulting in reduced sparsity. In contrast, C-MORL generates a dense and comprehensive Pareto front in these benchmarks. C-MORL also demonstrates the highest expected utility in nine of the ten benchmarks, which indicates that C-MORL can better maximize user utility given various user preferences. Appendix G.2 further illustrates that as a multi-policy approach, C-MORL can align corresponding Pareto-optimal solution given any preference. This is in contrast to single preference-conditioned policies, which may yield dominated solutions for unseen preferences.
Regarding time complexity, as stated in Section 4, C-MORL is linear time complexity with respect to the number of objectives, and is therefore generalizable to benchmarks with more than three objectives. We also note from Table 2 that for the Building-9d benchmark with nine objectives, the training times for PG-MORL and GPI-LS are excessively long, exceeding the time limitation we set. This finding aligns with the time complexity analysis of PG-MORL discussed in Section 4. Additionally, in both Pareto initialization and extension stage, the policies can be trained in parallel, which further reduces the total running time. Given that the expected running time is $O(nKN)$, the Pareto extension steps $K$ and number of selected policies $N$ can be adjusted to balance performance with running time.
Environments Metrics CAPQL Q-Pensieve PG-MORL GPI-LS C-MORL
MO-Hopper-2d HV$(10^{5})$ 1.15$\pm$0.08 1.26$\pm$0.01 1.20$\pm$0.09 1.19$\pm$0.10 1.39$\pm$0.01
EU$(10^{2})$ 2.28$\pm$0.07 2.28$\pm$0.01 2.34$\pm$0.10 2.33$\pm$0.10 2.56$\pm$0.02
SP$(10^{2})$ 0.46$\pm$0.10 1.61$\pm$1.31 5.13$\pm$5.81 0.49$\pm$0.37 0.33$\pm$0.28
MO-Hopper-3d HV$(10^{7})$ 1.65$\pm$0.45 1.66$\pm$1.20 1.59$\pm$0.45 1.70$\pm$0.29 2.29$\pm$0.01
EU$(10^{2})$ 1.53$\pm$0.28 1.26$\pm$0.79 1.47$\pm$0.25 1.62$\pm$0.10 1.80$\pm$0.01
SP$(10^{2})$ 2.31$\pm$3.16 1.77$\pm$0.88 0.76$\pm$0.91 0.74$\pm$1.22 0.28$\pm$0.09
MO-Ant-2d HV$(10^{5})$ 1.11$\pm$0.69 2.55$\pm$0.54 0.35$\pm$0.08 3.10$\pm$0.25 3.13$\pm$0.20
EU$(10^{2})$ 2.16$\pm$0.94 3.14$\pm$0.49 0.81$\pm$0.23 4.28$\pm$0.19 4.29$\pm$0.19
SP$(10^{3})$ 0.18$\pm$0.07 3.63$\pm$2.71 2.20$\pm$3.48 3.61$\pm$2.13 1.67$\pm$0.85
MO-Ant-3d HV$(10^{7})$ 1.22$\pm$0.33 3.82$\pm$0.43 0.94$\pm$0.12 0.55$\pm$0.81 4.09$\pm$0.13
EU$(10^{2})$ 1.30$\pm$0.29 2.18$\pm$0.41 1.07$\pm$0.07 2.41$\pm$0.20 2.57$\pm$0.06
SP$(10^{3})$ 0.17$\pm$0.09 0.83$\pm$0.07 0.02$\pm$0.01 1.96$\pm$0.79 0.03$\pm$0.01
MO-Humanoid-2d HV$(10^{5})$ 3.30$\pm$0.05 0.90$\pm$0.62 2.62$\pm$0.32 1.98$\pm$0.02 3.43$\pm$0.06
EU$(10^{2})$ 4.75$\pm$0.04 1.89$\pm$0.45 4.06$\pm$0.32 3.67$\pm$0.02 4.78$\pm$0.05
SP$(10^{4})$ 0$\pm$0.00 1.08$\pm$1.39 0.13$\pm$0.17 0$\pm$0.00 0.18$\pm$0.27
Building-3d HV$(10^{12})$ 0.33$\pm$0.18 1.00$\pm$0.02 0.83$\pm$0.02 0.26$\pm$0.04 1.15$\pm$0.00
EU$(10^{4})$ 0.75$\pm$0.09 0.96$\pm$0.00 0.93$\pm$0.01 0.74$\pm$0.01 1.02$\pm$0.00
SP$(10^{4})$ 1.83$\pm$0.75 9.23$\pm$7.77 0.37$\pm$0.22 0.67$\pm$0.90 0.69$\pm$0.62
Building-9d HV$(10^{31})$ 4.29$\pm$0.73 7.28$\pm$0.57 T/O T/O 7.93$\pm$0.07
EU$(10^{3})$ 3.31$\pm$0.06 3.46$\pm$0.03 T/O T/O 3.5$\pm$0.00
SP$(10^{3})$ 4.34$\pm$3.72 1.04$\pm$0.38 T/O T/O 2.79$\pm$0.40

### v2

# Efficient Discovery of Pareto Front for Multi-Objective Reinforcement Learning

@@ Abstract
Multi-objective reinforcement learning (MORL) excels at handling rapidly changing preferences in tasks that involve multiple criteria, even for unseen preferences. However, previous dominating MORL methods typically generate a fixed policy set or preference-conditioned policy through multiple training iterations exclusively for sampled preference vectors, and cannot ensure the efficient discovery of the Pareto front. Furthermore, integrating preferences into the input of policy or value functions presents scalability challenges, in particular as the dimension of the state and preference space grow, which can complicate the learning process and hinder the algorithm’s performance on more complex tasks. To address these issues, we propose a two-stage Pareto front discovery algorithm called Constrained MORL (C-MORL), which serves as a seamless bridge between constrained policy optimization and MORL. Concretely, a set of policies are trained in parallel in the initialization stage, with each optimized towards its individual preference over the multiple objectives. Then, to fill the remaining vacancies in the Pareto front, the constrained optimization steps are employed to maximize one objective while constraining the other objectives to exceed a predefined threshold. Empirically, compared to recent advancements in MORL methods, our algorithm achieves more consistent and superior performances in terms of hypervolume, expected utility, and sparsity on both discrete and continuous control tasks, especially with numerous objectives (up to nine objectives in our experiments). Our code is available at https://github.com/RuohLiuq/C-MORL.
@@ 1 Introduction
In many real-world control and planning problems, multiple and sometimes even conflicting objectives are getting involved. Such situations necessitate striking a better trade-off among these decision-making goals (Roijers et al., 2013; Hayes et al., 2022). For instance, in industrial control scenarios (Salvendy, 2001; Wang et al., 2023), maximizing utility and minimizing energy consumption are of particular interest as objectives to be optimized. Since different decision makers have heterogeneous preferences over these objectives, there may exist multiple Pareto-optimal policies (Roijers et al., 2014). Classical reinforcement learning (RL) methods typically involve training individual policies exclusively to align with each preference weight vector over multiple rewards (Nagabandi et al., 2018; Gupta et al., 2018). Yet it may lead to an enormous computational burden due to the overly dependence on the model retraining and fine-tuning stages. Moreover, such policies are hard to directly generalize or transfer to newer tasks (Cobbe et al., 2019; Taiga et al., 2022). Therefore, the multi-objective reinforcement learning (MORL) paradigm has drawn significant attention by reformulating these tasks for optimizing towards multiple criterion (Xu et al., 2020; Basaklar et al., 2022; Zhu et al., 2023). MORL aims to obtain either a single policy readily adapted to different preferences (Felten et al., 2023b; 2022; Teh et al., 2017) or a set of policies (Zhao & Grover, 2024; Felten et al., 2024; Röpke et al., 2024; Kim et al., 2024) aligned with their respective preferences.
@@
Our algorithm achieves favorable time complexity characteristics and derives a high-quality Pareto front, as indicated by the following observations. Firstly, training several initial policies based on fixed preferences is efficient. In addition, the adoption of constrained update steps allows for rapid adjustment of the initial policy, leading to the derivation of new solutions on the Pareto front. Regarding the second research gap concerning the complete Pareto front, different from the meta-policy approach that relies on a single initial policy (Chen et al., 2019), our method can extend multiple policies selected from initial policies based on their crowd distance to enhance diversity while promoting better performances. Intuitively, a larger crowd distance indicates that the corresponding policy appears on a sparser area on the Pareto front, therefore, extending such a policy is more likely to fill the Pareto front. While sharing similarities with our formulation, as one category of multi-objective optimization (MOO) approach, epsilon-constraint methods solves a MOO problem by converting it to several single-objective constrained optimization problems (Laumanns et al., 2006; Van Moffaert & Nowé, 2014). However, the running time of such a method is exponential in problem size, rendering it impractical for MOO problems with numerous objectives. In contrast, crowd-distance-based policy selection eases such burden, exhibiting linear complexity, and can efficiently solve MORL tasks with numerous objectives. Our main contributions are as follows:
@@ 2 Related Work
Prior trials on tackling multi-objective RL fall into training single preference-conditioned policy or multi-policy. Typical single-preference-conditioned policy approaches adopt a policy that takes preferences as part of network inputs and utilizes the weighted-sum scalarization of the value functions (or advantage functions) to optimize the policy  (Van Moffaert et al., 2013; Parisi et al., 2016; Yang et al., 2019; Zhang & Golovin, 2020; Basaklar et al., 2022; Lu et al., 2022; Hung et al., 2022; Zhu et al., 2023; Lin et al., 2024). In the evaluation stage, agents can execute corresponding solutions based on users’ desired preferences (Yang et al., 2019; Basaklar et al., 2022). With the shared neural networks, gradients from different tasks can interfere negatively, making learning unstable, leading to imbalanced performance across the entire preference space, and sometimes even less data efficient.
@@
Instead of training a single policy agent, multi-policy approaches train a finite set of policies to approximate the Pareto front (Abels et al., 2019; Xu et al., 2020; Alegre et al., 2022). In the evaluation stage, the policy in policy set that maximizes the utility of the user (i.e., the weighted sum of objectives) is chosen (Friedman & Fontaine, 2018). Rather than evenly sampling from preference space, (Xu et al., 2020) proposes an efficient evolutionary learning algorithm to find the Pareto set approximation along with a prediction model for forecasting the expected improvement along vector-valued objectives. Yet the performance highly depends on prediction model’s accuracy, and it is challenging to recover the Pareto front due to the long-term local minima issue. This method also suffers from low training efficiency due to the exponential complexity involved in repeatedly calculating the quality of the virtual Pareto front during the training process. (Kyriakis & Deshmukh, 2022) designs a policy gradient solver to search for a direction that is simultaneously an ascent direction for all objectives. (Alegre et al., 2022) firstly trains a set of policies whose successor features form a $\epsilon$-$CCS$ (convex coverage set), then utilizes the generalized policy improvement (GPI) algorithm to derive a solution for a new preference. (Alegre et al., 2023) further introduces a novel Dyna-style MORL method that significantly improves sample efficiency while relying on accurate learning of dynamics. (Röpke et al., 2024) discovers the Pareto front by bounding the search space that could contain value vectors corresponding to Pareto optimal policies using an approximate Pareto oracle, and iteratively removing sections from this space. Our proposed C-MORL distinguishes from their divide-and-conquer scheme by a novel policy selection procedure along with explicitly solving a principled constrained optimization.
@@
Our approach bears connections with (Chen et al., 2019) and (Huang et al., 2022), while distinguishing itself in terms of both the training process and objectives. Compared to the meta-policy approach, adaptation process is not necessary in the evaluation phase of our method. When presented with an unseen preference, the policy in the Pareto front with the highest utility is chosen as the surrogate execution policy. In contrast to (Huang et al., 2022; Kim et al., 2024), we do not aim at solving a constrained RL problem for training the working policy under specific preference. Rather, we propose to leverage the constrained optimization steps to fill the complete Pareto front with enhanced flexibility (Liu et al., 2020; Xu et al., 2021), and design extension and selection algorithms to explicitly promote diverse policies. Constrained optimization techniques are widely adopted to solve multi-objective optimization (MOO) problems. Epsilon-constraint methods are a category of MOO techniques that involve pre-defining a virtual grid in the objective space and solving single-objective problems for each grid cell, where the optimum of each problem corresponds to a Pareto-optimal solution (Laumanns et al., 2006). Instead of pre-defining virtual grid, this work proposes crowd distance based policy selection to address the exponential complexity in the epsilon-constraint method.
@@ 3.1 MORL Setup
The goal of MORL is to find a policy so that each objective’s expected return in $\bm{G}^{\pi}$ can be optimized. In practice, since training RL typically requires a scalar reward to interact with the training agent, and co-optimizing multiple objectives is hard to achieve an ideal tradeoff, especially in a situation where objectives are competing against each other. To that end, denote $\Omega=\{\bm{\omega}\in\Omega|\sum_{i=1}^{n}\omega_{i}=1,\omega_{i}\geq 0\}$ as the preference vector. We use the preference function $f_{\bm{\omega}}(\mathbf{r})$ to map a reward vector $\mathbf{r}(\mathbf{s},\mathbf{a})$ to a scalar utility given $\bm{\omega}:f_{\bm{\omega}}(\mathbf{r}(\mathbf{s},\mathbf{a}))=\bm{\omega}^{\top}\mathbf{r}(\mathbf{s},\mathbf{a})$. Our goal is then to find a multi-objective policy $\pi(\mathbf{a}|\mathbf{s},\bm{\omega})$ such that the expected scalarized return $\bm{\omega}^{\top}\bm{G}^{\pi}$ is maximized.
@@ 3.2 Pareto Optimality
Figure 1: Visualization of metrics. (a) Hypervolume, reference point, and example of crowd distance calculation. As an example, the crowd distance of $\pi_{b}$ is calculated based on the expected return of its neighbors $\pi_{a}$ and $\pi_{c}$, as well as the extreme solutions on the two objectives. (b) Given a preference vector, the corresponding expected return is calculated by selecting the set max policy from Pareto front solutions.
@@ 4 Optimizing Constrained MORL
Assumption 4.2 justifies the use of gradient-based algorithm for solving CMDP can converge with general constraints. As long as the threshold $\mathbf{d}$ are chosen beforehand such that for initializations, there are $G_{i}^{\pi}\geq d_{i},i\neq l$, then starting points for problem Eq. 1 is always feasible.
@@
$D^{*}\triangleq\min_{\bm{\lambda}}\max_{\pi}\;L(\bm{\lambda},\pi).$ (2)
@@ 5.3 Pareto Extension
where $G_{i}^{\pi_{r}}$ is the expected return of the $i^{th}$ objective in the last constrained optimization step, which is indexed by $r$ for the current iteration. Next, we introduce practical methods to solve Eq. 3.
C-MORL-IPO Constrained Policy Optimization (CPO) is a widely used general-purpose policy search algorithm for constrained RL (Achiam et al., 2017). While the CPO algorithm ensures monotonic policy improvement and guarantees constraint satisfaction throughout the training process, inner-loop optimization is required when there are multiple constraints. Therefore, CPO is hard to handle more than two MORL objectives. To overcome this issue, inspired by (Xu et al., 2021; Liu et al., 2020), we also propose to find a satisfiable solution efficiently for Eq. 3 via the interior point method (IPO), which is a primal type approach and holds the promise of finding a solution closer to the original C-MORL as shown in Theorem 4.5. Given $G_{i}^{\pi_{r}}$ and the the defined log barrier in 4.5, we then convert problem Eq. 1 into an unconstrained optimization problem:
@@
The detailed procedures of solving Eq.4 are provided in Appendix F. In practice, IPO is more robust in stochastic environments, and larger $t$ would guide to a solution with higher rewards yet with more computation costs. Without the requirement of calculating the second-order derivative, C-MORL-IPO is more computationally efficient than trust-region-based method for CPO updates (Achiam et al., 2017), which we term as C-MORL-CPO and compare it in Appendix F.
@@ 6 Experiments
In this Section, we validate the design of our proposed algorithm using both popular discrete and continuous MORL benchmarks from MO-Gymnasium (Felten et al., 2023a) and SustainGym (Yeh et al., 2024). These benchmarks include five comprehensive domains: (i) Grid World includes Fruit-Tree, a discrete benchmark with six objectives. (ii) Classic Control includes MO-Lunar-Lander, a discrete benchmark with four objectives. (iii) Miscellaneous includes Minecart, a discrete benchmark with four objectives. (iv) Robotics Control includes five MuJoCo tasks with continuous action space based on MuJoCo simulator (Todorov et al., 2012; Xu et al., 2020). (v) Sustainable Energy Systems includes two building heating supply tasks. These benchmarks pose significant challenges for MORL due to complex state (up to $348$)/action (up to $23$) spaces and large objectives spaces (up to $9$ objectives). More details of experiment settings are listed in Appendix E.
We benchmark our algorithm against five competitive baselines under the metrics of hypervolume (HV), expected utility (EU), and sparsity metrics (SP). Higher HV and EU, lower SP indicate better performance. Each of the baselines are trained for $5\times 10^{5}$ time steps for discrete benchmarks. Continuous benchmarks with two, three, and nine objectives are trained for $1.5\times 10^{6}$, $2\times 10^{6}$, and $2.5\times 10^{6}$ steps, respectively. The baselines involve: (i)Envelope (Yang et al., 2019).(ii) CAPQL (Lu et al., 2022). (iii)Q-Pensieve (Hung et al., 2022). (iv) PG-MORL (Xu et al., 2020). (v) GPI-LS (Alegre et al., 2023). (vi) MORL/D (Felten et al., 2024). Among these baselines, Envelope is developed for discrete control tasks, GPI-LS is suitable for both discrete and continuous control. The other baselines are specifically tailored for continuous control.
@@
While C-MORL leads the highest hypervolume by a large margin in almost all cases, it does not always attain the best sparsity in some benchmarks. This occurs because, in certain cases, an algorithm may identify only a few similar Pareto-optimal solutions, leading to low sparsity. As illustrated in Table 2, along with Figures 4 (c) and 5, both CAPQL and GPI-LS cover only a specific portion of the Pareto front, resulting in a reduced sparsity. In contrast, C-MORL generates a dense and comprehensive Pareto front in these benchmarks. C-MORL also demonstrates the highest expected utility in nine of the ten benchmarks, which indicates that C-MORL can better maximize user utility given various user preferences. Appendix G.3 further illustrates that as a multi-policy approach, C-MORL can align corresponding Pareto-optimal solution given any preference. This is in contrast to single preference-conditioned policies, which may yield dominated solutions for unseen preferences.
Regarding time complexity, as stated in Section 4, C-MORL is linear time complexity with respect to the number of objectives, and is therefore generalizable to benchmarks with more than three objectives. We also note from Table 2 that for the Building-9d benchmark with nine objectives, the training times for PG-MORL and GPI-LS are excessively long, exceeding the time limitation we set. This finding aligns with the time complexity analysis of PG-MORL discussed in Section 4. Additionally, in both the Pareto initialization and extension stages, the policies can be trained in parallel, which further reduces the total running time. Given that the expected running time is $O(nKN)$, the Pareto extension steps $K$ and the number of selected policies $N$ can be adjusted to balance performance with running time.